Sort by:[Date]

4-tile hole in <ADC> and <TAC> QA plots

 In spring 2018, we were seeing occasional runs in which a group of four tiles would have low <ADC> and <TAC>.

Totally dead tile

 Starting in run 20127001 the tiles associated with West PP2 TT04, linked to channel EQ1 Bd0x1D Ch 10 was dead.

status message density plot

EPD Cold tiles again

Executive Summary

analysis of eTOF data before / after HV change

Attached are the matching ratio plots for several runs before the HV change combined ( mostly day345) and after the HV change (mostly day 347).

EPD cable weight between Rx boards and QT boards

The take away: regularly check the cable tension between Rx boards and QT boards!

EPD channel measuring cosmics out of Rx board, but blue in QA plots and not showing any MIP peaks

 At the beginning of Run 20 (December 2019), we noticed that a single EPD tile (West PP3 TT24) was showing no signs of life.

Holes in EPD QA plots

 In run 19, we were seeing that, over the course of a few hours, the <ADC> displayed in the QA plots for some of the tiles would slowly become blue.

Y joints breaking

 This page is very brief. There are Y-shaped joints in the fiber bundles, where the bundle splits in two.

Overview of Recent Heavy-Flavor and Jet Results from the STAR Experiment